OpenStudy (anonymous):

Let f(x) = x + 2 and g(x) = x2 – 6x + 3. Find g(f(x))

OpenStudy (anonymous):

in g(x) replace 'x' with 'x+2' and simplify

OpenStudy (anonymous):

well im not simplifying right cause i keep getting it wrong

OpenStudy (ghazi):

\[g(f(x))=(x+2)^2-6(x+2)+3\] i hope you can do it now :)

OpenStudy (anonymous):

i did that part but i keep getting x^2-4x-9 as my answer which is wrong

OpenStudy (anonymous):

you should get x^2-2x-5

OpenStudy (anonymous):

i know but i dont know how to get there

OpenStudy (anonymous):

okay. First expand (x+2)^2. what do you get?

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

ohhh i see what i did wrong thank you

OpenStudy (ghazi):

\[g(f(x))=(x^2+2x+4)-6x-12+3=x^2-4x-5\]

OpenStudy (anonymous):

very good.